Safety and Database Subsection ADO - PO - SA Main safety projects / activities RP issues Patrols What is the current status and what needs to be implemented before the end of the long shutdown 12.03.2009 , Olga Beltramello. General Safety
Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.
Safety and Database Subsection
ADO - PO - SA
A new safety organization have been implemented at Point 1with :
The SLIMOS desk has been set up operational for the LHC run period and later moved to shutdown configuration.
All safety systems and access systems (LACS/LASS, SSA) IHM are implemented and operational.
The SLIMOS procedures are available.
The SLIMOS handbook is created and under consolidation, this shall be finalized before the end of the shutdown period.
The Fire Brigade instructions and routes (with the associated signalization in situ) have been implemented (Sofia, Irina and Kathy).
The safety systems (ODH, FG, Fire, seismograph), access systems and some missing infrastructure (ventilation) data are currently under collection and shall be inserted in the SLIMOS desk and also in the safety web page (Alessandro).
The SLIMOS shifts are operational, about 50 SLIMOS have been trained and qualified (Sofia and Xavier).
Safety Systems (Fernando, Sofia, Nuno)
Safety systems / safety matrix status :
An ATLAS training for Fire Brigade has been developed (Sofia, Kathy and Xavier) and the entire fire brigade has been for half a day trained.
Very appreciated and successful !
An ATLAS training for visitor guides has been developed to cope with the new access and RP situation (more than 50 people trained by Xavier).
The ATLAS Safety training is currently under revision :
Main Safety Engineering consolidation campaigns before this summer
A campaign to improve the overpressure between the caverns and
the tunnel is currently on going in collaboration with CV (air flow
measurements, sealing improvement, overpressure measurement,
implementation of additional pressure or air flow sensors, link
to the slimos desk, etc..) (Olga, Fernando, Nuno)
An assessment of the criticality of possible CO2 leaks during ID
operation is being performed with the various TRT and ID experts.
with the study of the possible technical and operational solutions.
A large CO2 tests campaign will be run in collaboration with ID
experts and Fire Brigade to assess the stability of the various CO2
systems, the various operation parameters and the safety
RP issues – General
The RPE / RPA team has been trained, the piquet service has been set up and the ATLAS RP activities have been successfully organized and run since the first beams (Giuseppe).
Additional training is possible with N. Conan from SC-RP every week , please register.
A new session of RPA and RPE training is planned this year, if you are interested please contact Giuseppe or Olga.
The ATLAS activation and dose calculations is currently under restructuring and revision (Giuseppe and Zusanna)
RP issues – Gate monitors
The Radiation gate monitors to detect possible radioactive material exiting at UX15 access points have been selected in collaboration with SC-RP and ordered to Dubna (available end of March).
The infrastructure and access points configuration need to be modified to implement the gate monitors, currently under study (Tatiana and Wim).
RP issues – Surface Buffer Zone
The creation of the surface radioactive buffer zone in the surface building is under study (Patrick) . It shall be implemented before the next shutdown period.
RP issues – ALARA
In order to gain experience when the dose rates are negligible and to prepare for future shutdowns, we have agreed with Michel to follow the coming ATLAS Detector closure and to collect the data required for the ALARA studies.
The strategy is the following :
RP issues – ALARA
RP issues – Upgrade studies
We have started our participation to the Upgrade studies in collaboration with Michel for the Lar Upgrade and Raphael for the IBL and beam pipe replacement :
The ATLAS patrols procedures are ready and tested and
The Patrols are under the responsibility of the GLIMOS.
The ATLAS patrol manager is William.
Several ATLAS patrols have been run with around 30 persons from ATLAS Collaboration and DT persons.
The patrollers and head of patrol training and validation have to be created and organised.
The patrol team piquet on call service are currently under setting up ( for the LHC run period)
This on call service shall be associated to other services like RPE/RPA to minimize impact on personnel.
No more than 3 to 5 patrols should be expected per year.